EVA Report – February 2nd

Crew 188 EVA #5 Report 2FEB2018

Author’s name: Tatsunari Tomiyama AHFP, Health & Safety Officer

Purpose of EVA: To perform waypoint data collection of North to Galileo road and scout for future EVA sights

Location of EVA: Galileo Road, and Cactus Road

UDM27 Coordinates: 518800 E, 4250600 N and 519200 E, 4252000 N

Number of EVA Crew: 3

Participants: Julia De Marines, *Sarah Jane Pell, and Tatsunari Tomiyama

*Sarah Jane subbed in for Renee due to assessing her fatigue level in the morning and determining she needed to rest and postpone her EVA to tomorrow.

EVA Commander: Tatsunari Tomiyama

Rovers used: Spirit and Curiosity

Duration: 2 hours

EVA Departure Time: 10:50am

EVA Return Time: 12:45pm

Summary:

We were delayed in our start time for this EVA due to having to re-suit Julia in an alternate EVA suit due to suit #2 having a broken switch. The EVA objective was to collect waypoint data for future EVA sights. This EVA expected and planned to collect waypoints for Galileo Road, and several points on the Cactus Road. Overall, we successfully completed the waypoint data collection on the Cactus Road and Galileo Road.

There were 4 check points on the Cactus Road on the planning. However, we made 6 waypoints on the data collection in this EVA due to the clear visualization. We spent more time than we planned to return to the Hab due to the sample collection being performed at the waypoints.

We made 4 waypoints during this EVA. One is at Galileo Road with the Cow Dung Road as we were trained on driving ATV and rovers. Second waypoint was the intersection with Cactus Road on Galileo Road. This waypoint was relatively difficult to recognize the road visually. We went to the East on cactus Road first and made a waypoint at 11:04. Then, we turned back to the north on the Cactus Road. We stopped and made a waypoint at the entrance of Candor Chasma on Cactus road. We kept driving on Cactus Road to the north. We made a final planned waypoint to the Yellow Zebra Rd. Finally, we visited to the north end of the Cactus Road before reversing our route and returning to the Hab.

EVA Report – February 1st

Crew 188 EVA #4 Summary Report 1FEB2018

Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander

Purpose of EVA: To test human endurance on a pedestrian EVA to Phobos Peak and test the 360 video capture in the field.

Location of EVA: Phobos Peak Rock
UDM27 Coordinates: 519260 E, 4250260 N
Number of EVA Crew: 3
Participants: Julia DeMarines (EVA CDR), Sarah Jane Pell, and Ryan Kobrick EVA Commander: Julia DeMarines Rovers used: none

Duration: 4 hours 14 minutes
EVA Departure Time: 10:44am
EVA Return Time: 14:12pm

Summary:

The EVA had three primary objectives that were all completed today by the crew:
1 – sample potential “green” samples for Julia’s chlorophyll detection project with NASA Ames
2 – shoot Insta-360 Pro high definition footage for Sarah Jane’s Monash Immersive Visualization Platform project
3 – summit Phobos peak while collecting biometric data for Ryan’s EVA metrics study funded by Florida Space Research Program

The crew faced equipment challenges (see Sol 4 Summary Report) but was able to overcome and persevere to meet the mission goals. The investigators have high hopes for the data collected, but regardless experienced a true Martian adventure.

The crew walked to Phobos Peak with two stops to set up the 360 camera, one stop for sampling, and several stops for interesting formations. We summitted the mountain by scrambling up the West face (the side you can see form the Hab) with several safety stops to combat fogging and shoe repair issues. The international trio waved their 3 flags from the summit for a mini photoshoot from the Hab. On the walk back from the successful mission, the crew stopped for a rare opportunity to toss around a bouncy ball on Mars testing their reflexes, spacesuit glove dexterity and coordination. A radio was dropped near the end of the EVA, but using the GPS track the crew quickly recovered the valuable asset. [end]

EVA Report – January 30th

Crew 188 EVA #3 Summary Report 30 January 2018
EVA Date: 30 Jan 2018
Author: Tatsunari Tomiyama

Purpose of EVA: Ground truthing to establish way points for more accurate future missions, as well as to capture “B roll” of the terrain and us walking for use in video productions. A tertiary goal is to put hours on the rover batteries.

Location of EVA: Lith Canyon
UDM27 Coordinates: 518800 E, 4256500 N
Number of EVA Crew: 3
Crew Members Going on EVA: Ryan Kobrick (Commander), Zac Trolley (Crew Engineer) and Tatsunari Tomiyama (Health & Safety Officer)
EVA Commander: Tatsunari Tomiyama
Rovers used: SPIRIT and CURIOSITY
Rover usage: 2
EVA Duration: 3 hour 27 minutes
EVA Departure Time: 10:24am
EVA Return Time: 13:15pm

Time Check:
09:48 Prep Start + Simulation Start
10:16 Airlock
10:21 EVA Start
10:24 ATV Leave
12:30 Arrived Parking Lot
12:41 Additional near-hab duties, photos and videos
13:01 Airlock
13:15 EVA End

Summary: Our ground truthing reconnaissance was planned to explore the “Goblins” near Lith Canyon. However, we had some difficulty finding our GPS waypoints goals and spent a little more time than planned looking for checkpoints. The planned EVA route was to reach the “Goblins” by way of Cow Dung Road passing Pooh’s Corner and all the way to Lith Canyon Rd 1106. We reached the end of Lith Canyon Rd 1106, took photo and video documentation and began our return to the Hab. Additionally, data was collected for Dr. Kobrick’s EVA for a metrics study on human performance physiology and exploration metrics. We set up checkpoints on this EVA at the following locations: Cowboy Corner, Tank Wash, Intersection to Brahe Hwy 1572, intersection to the Quarry Rd, and the intersection to the Lith Canyon 1106.
On the return trip to the Hab, we collected samples of Gypsum between the intersection to the Brahe Hwy 1572 on the Pooh’s Corner and Tank Wash. We will consult with our crew scientist and mission support team to analyze the sample in the science lab. Once the rovers were parked, we walked around the Hab to get more photos before heading back into the airlock.

EVA Report – January 29th

EVA #2

EVA Date: 29 Jan 2018

Crew Executive Officer Reporting

Purpose of EVA: Spacesuit activity recognition and assessment of in-suit performance
Location of EVA: South of Marble Ritual (walking), Pooh’s corner (vehicle)
UDM27 Coordinates: 518800 E, 4250600 N

Number of EVA Crew: 3

Crew Members going on EVA:
Renee Garifi (Executive Officer), Julia De Marines (Crew Astronomer), Zac Trolley (Crew Engineer)

Rovers used:
– DEIMOS
– Opportunity
– Spirit

Rover time usage: 0.5 hours

Duration of EVA: 1.3 hours

EVA departure time: 15:10
EVA return time: 16:42

Time Checks:
Start Prep: 15:32
Airlock Start: 15:10
Roll out: 15:20
Check-In: 15:32
Comms on Overshoot: 15:40
Scrub due to helmet issue: 15:54
Renee in Airlock: 15:58
Zac and Jules return to get Rover: 15:59
Jules in Airlock: 16:15
Zac in Airlock 16:28
Equipment Cleaned and put away: 4:42

Summary: The Crew attempted to locate the Marble Ceremony landmark near the Pooh’s Corner rock area. They attempted to utilize one of the hab GPS units to identify the planned EVA coordinates. We took with us one hab GPS unit, as well as one of the small, laminated MDRS, printed maps to find our waypoints. Our objective was to perform a training EVA to familiarize our crew (in two team groups of three) with EVA operations first-hand from a participant perspective. All members brought cameras for photos and video footage to send to Mission Support. All three rovers used were fully functional and tested in the rough paths of the Martian Terrain on Cow Dung Road.

We experienced an off-nominal situation with one crewmember during this second EVA of the day. All three of the crew experienced a higher-than-normal amount of helmet fogging while walking. It was realized shortly after parking the rovers on cow dung road that we had overshot the EVA destination and were too far north. While walking back to the rovers, one crewmember experienced helmet fog to the point of having impaired visibility and was unable to drive their rover. After waiting a few minutes for the fogging to clear, the crewmember began to have sweat drip down into the eyes and collect in the glasses. They requested assistance from another crewmember to turn the rover around for them. In the process, the crewmember having the fogging issue experienced a total visual loss when their hat fell from their head and into the front of their helmet, completely blocking their view.

When this hat anomaly occurred, the crewmember made the call to be driven back to the hab to receive assistance with the helmet from the HabCom team. All three crewmembers returned on two rovers, dropping the impaired crewmember at the airlock and returning to retrieve the third rover.

The remaining two crewmembers received permission to resume the EVA and made it out to Pooh’s corner safely and completed the marble ceremony. Despite having to return one crewmember to the hab early, this EVA was highly successful for three reasons. First, it provided a non-life threatening emergency that required immediate assistance for the crew to respond to.

There is a long list of reasons why an EVA will terminate early, so we are fortunate that this instance was minor and could be resolved by removing the helmet back at the hab. Secondly, the crew of three proved they could work together in an off-nominal situation to problem solve and quickly make a decision that places the safety of the crew above the objective of the EVA. Lastly, it allowed us to demonstrate our contingency plan for returning to the hab if one of the rovers breaks down since one rover had to remain parked while all crew returned to the hab.

The crew engineer extended their EVA in order to service the ATVs and Rovers that we used during the day.

EVA Report – January 29th

EVA #1

EVA Date: 29Jan2018

Health & Safety Officer Report

Purpose of EVA: Development of Situational Awareness for Spacesuit Activity

Location of EVA: South of Marble Ritual (walking), Pooh’s corner (vehicle)

UDM27 Coordinates: 518800 E, 4250600 N

Number of EVA Crew: 3

Crew Members Going on EVA: Ryan Kobrick (Commander), Sarah Jane Pell (Residence-in-Artist), Tatsunari Tomiyama (Health & Safety Officer)

ATV used: 1,2,3

ATV usage: 5 minutes

EVA Duration: 1 hour 7 minutes

EVA Departure Time: 13:30

EVA Return Time: 14:32

Time Check:

13:01 Prep Start + Simulation Start
13:25 Airlock
13:30 EVA Start
13:35 ATV Leave
13:40 Parked
14:02 Arrived Ritual Marbles
14:18 Return to Hub
14:27 Arrived Parking Lot
14:32 Airlock

Summary: Crews attempted EVA to develop situational awareness for space suit performance. We used 3 ATVs to travel to the targeting location; Marble Rituals. When we arrived at the nearest place on Pooh’s corner, we parked ATV and started to walk. We found a fossil and recent animal footprint soon. Once arrived, we take photos and investigated the location. Then, we returned to the Hub. Overall, there was no significant issue to this EVA and we enjoyed the first simulation.

EVA Report – January 24th

EVA Report for 24Jan2018

Author: Oscar Ojeda

Purpose of EVA: Usage of rovers for battery discharge, aerial mapping of terrain, testing of mid-range EVA protocols, testing of the Guache I Rover, Collection of possible halophile samples.

Destination: Intersection of Quarry Road and Cow Dung Road

Coordinates: 518400 E, 4255500 N

Participants: Atila, Danton, David and Oscar

Narrative: One successful EVA was undertaken on this day. A crew of 4 departed the habitat at 11:35, aiming for the two main goals of discharging the rover’s battery and to keep testing the different projects. After taking Cow Dung Rd for about 30 minutes, the crew arrived at Quarry Road, where several aerial shots were taken, testing the 3D mapping software, as well as a first field run of the Guache I rover. On the way back, the crew stopped on two more locations to perform the same activities, taking aerial footage, and testing the Guache I rover performance over different types of terrain. The expedition never left the main road for more than 20 meters. No halophile samples were taken, because no relevant sites were found during the expedition. The batteries of the rovers were discharged, and in particular, opportunity dropped to 55% during the traverse. The crew entered the airlock at 13:55 and finished the EVA.

No incident or anomaly was sustained during the EVA.

EVA Report – January 25th

EVA Report for 25Jan2018

Author: Oscar Ojeda

Start time: 10:00

End time: 15:00

Purpose of EVA: Aerial mapping of terrain, testing of long-range EVA protocols, testing of temperature regulation vest, Collection of possible halophile samples.

Destination: Candor Chasma

Coordinates: 520500 E, 4251000 N

Participants: Atila, Luis and David and Tania

Narrative: The exploration team left the Hab at 10:35. Being the last exploration for the majority of the crew, they were willing to go to the depths of Candor Chasma and conquer it. They took the 4 ATVs and advanced on the main route until arriving at Galileo Road 1104 with the intention of entering the canyon from an area near the intersection between Stream Bed Connector and Cactus Road. Not being able to visualize Stream Bed Connector, the expedition team arrived Mountain Goat Trail and returned by the same way to find Stream Bed. Apparently, no one has used this path for a long time and it was impossible to find. Because of that, the expedition team returned to Galileo Road, and had a short communication with the Hab. They decided to advance approximately half a kilometer in the direction of Cactus Road and stop. They walked almost 10 minutes and found a perfect entrance for Candor Chasma. At 12:03 they entered into the depths of such majestic beast and advanced for 50 minutes. They found traces of Martian species that resemble terrestrial antelopes, but they had no encounter with any living species. At 12:55 they decided to take the road back and were ready to return home at 1:43 pm. The way back happened without incident and they arrived at the Hab 14:20.

No incident or anomaly was sustained during the EVA.

EVA Report – January 23rd

EVA Report for 23Jan2018

Author: Oscar Ojeda

Purpose of EVA: Aerial mapping of terrain, testing of long-range EVA protocols, and long-range operation test of Cóndor Space Suit Simulator, Collection of possible halophile samples.

Participants: Oscar, Cynthia, Tania, and Danton

Coordinates: 518500 E, 4256500 N

Narrative: Two EVAs were undertaken today, the second one consisted on driving through Cow Dung Road to discharge the battery od each of the new rovers, to increase their battery life. The crew departed at 16:00 and was returning at the hab at 17:00. One of the crewmember’s radio was drained during the return, so he took the center position in the formation, for safety.

The first EVA took a crew of 4 to Lith canyon. The goals of the EVA were attained. The crew left the airlock at 10:35, securing the elements to the ATVs and taking the specified route, taking Cow Dung road north. The crew arrived at the entrance of the Special Region and struggled for about half an hour to find the entrance to the rest of the road, which is not easily visible. After finding the entrance the crew proceeded with the rest of the vehicle road, taking Lith Canyon Road to it’s end. At this point, the ATVs were left and the crew started walking up a hill to have a better visual of the area and plan the next exploration steps. Aerial footage was also taken from the Drone. Walking exploration of Lith canyon was undertaken, walking west from the position where the ATVs were parked. The crew walked towards the western wall of the canyon, departing after into the eastern limit, where aerial footage was gathered. The zone presented a significant amount of cactus and hervibore presence evidence. The crew started the return walk by 14:05, arriving to the ATVs by 14:45. The ATVs were boarded, and the road back resumed. On the way back, a sample of soil containing salt deposits was taken for further analysis for Halophiles. The crew arrived to the hab at 15:35, ending the EVA.

EVA Report – January 22nd

This is the report for today’s EVA

Purpose of EVA: Primary: Long-range EVA protocol testing, Sample collection for possible Halophile identification, identification of relevant sampling sites, heating vest testing.

Secondary: Outreach and Educational graphic material collection. Relevant sample collection.

Participants: Atila, Danton, Luis, and David.

Coordinates: 517100 E, 4256400 N

Narrative: Today was the first long range EVA undertaken by the crew, and it was an eventful one, fortunately with a very happy ending. It is important to say that the temperature was a challenging element for the crew. The EVA started at 12:00 with the egress of the crew of four of the airlock. By pairs they boarded the two rovers and departed north by Cow Dung road, however, by the rapid discharge of the opportunity rover, the crew returned to the hab to rest, eat something, and take two ATVs instead of Opportunity. The EVA was resumed at 13:30 with a new egress of the crew, this time boarding Deimos and the 1 and 2 350 ATVs. The crew departed north until the north of Copernicus Hw, where they left the vehicles and started walking at 14:50. The crew walked about 15 minutes to a valley(38° 27.448N 110° 48.031W) and stayed there for about 5 minutes, where they took the route back, taking relevant samples in the way, as well as aerial footage(38° 27.0021N 110°48.565W). There were a lot of hervibore prints as well. While on Brahe Hwy the battery meter of Deimos suddenly started dropping fast, taking about 5 minutes to go from 71% to 26% charge, as reported by the Engineering officer. The crew decided to switch positions to reduce the weight, and kept going, however the charge kept dropping. The crew decided to walk along the vehicle, and afterwards to push it. The crew followed this procedure for 50 minutes. Afterwards they took the jumping cables (for battery charging) and pulled the rover with one of the ATVs until they were in sight of the crew in the hab. Unfortunately, the cables broke, and it was impossible to pull the rover further. At that point the crew was already 45 minutes above the closing time of the EVA window, and director Shannon went out to help them. The crew broke simulation, towed and pushed the rover up to the habitat and made an ingress to the hab, out of sim.

They were greeted with hot cocoa and blankets by the rest of the crew, who were happy to see them again.

EVA Report – January 21st

Purpose of EVA: Primary: Mid-range testing of the Cóndor Space Suit Simulator, Mid-range EVA protocol testing, Sample collection for possible Halophile identification, further exploration of El Dorado Canyon.

Secondary: Outreach and Educational graphic material collection. Relevant sample collection.

Participants: Oscar, Luis, David, and Tania.

Narrative: Today marked the last of the mid-range EVAs programmed, leading to longer and even more interesting exploration. The weather was favorable, but windy. The crew donned the suits and exited the airlock at 12:30, the four crewmembers then boarded the Rovers Spirit and Opportunity and followed the path planned, towards El Dorado canyon. After approximately 25 minutes of traverse, and after passing Robert’s Rock Garden, the Rovers were parked on the side of the road, and the expedition started walking with Southeast course, towards the entrance of the canyon, upon finishing the rocky plains, two crewmembers had to return to the Rover, given that the aerial drone was left behind. Upon reuniting, the crew resumed the traverse until the entrance of the canyon. More herbivore and cougar prints were found in the area, but not as fresh as on the previous EVA to the canyon. The crew entered the canyon and followed the dry riverbed. Along this path, water samples were taken, aerial footage with the drone, as well as traditional photography. The crew followed the path until the eastern exit of the canyon, heading south to meet Cow Dung road again. However, the sampling sites that were found during the firs EVA to the area were not found, since they probably belong to a subsidiary structure, south of El Dorado canyon. It is suggested that further detailed mapping of this area is undertaken, to identify all the morphologies present. While returning through Cow Dung road, heading north, a crewmember’s cell phone (used as a compass and camera) was lost, so a search formation was done, and the crew headed south, being able to find the device, and resuming the walk back north. The walk lasted about 25 minutes, reuniting with the Rovers and heading back to the habitat. The EVA ended at 15:20 with the ingress of the crew from the airlock to the habitat.

The aerial footage is being processed, as well as the performance of the Cóndor Space Suit Simulator, which will be subjected to detailed inspection tomorrow.

The only incident related with the EVA was the lost and recovery of the cell phone.